# The Convex Matching Distance in Multiparameter Persistence

### v1

@@ Abstract
We introduce the convex matching distance, a novel metric for comparing functions with values in the real plane. This metric measures the maximal bottleneck distance between the persistence diagrams associated with the convex combinations of the two function components. Similarly to the traditional matching distance, the convex matching distance aggregates the information provided by two real-valued components. However, whereas the matching distance depends on two parameters, the convex matching distance depends on only one, offering improved computational efficiency. We further show that the convex matching distance can be more discriminative than the traditional matching distance in certain cases, although the two metrics are generally not comparable. Moreover, we prove that the convex matching distance is stable and characterize the coefficients of the convex combination at which it is attained. Finally, we demonstrate that this new aggregation framework benefits from the computational advantages provided by the Pareto grid, a collection of curves in the plane whose points lie in the image of the Pareto critical set associated with functions assuming values on the real plane.
@@ 1 Introduction
In recent years, research on multiparameter persistent homology has grown substantially, driven by the need to apply Topological Data Analysis (TDA) techniques to data represented by multifiltrations, as occurs, for example, when analysing or comparing data described by vector-valued functions [FrMu99, BiCeFrGiLa08, CaZo09, CaDFFe10, CaSiZo10, CaLa11, CaEtFrKaLa13, CeLa13, CaFeLa15, CeFr15, Le15, AlKaLa17, HaOtScTi19]. However, a fully developed theoretical framework enabling efficient applications of TDA in this setting is still lacking. A widely adopted comparison method in multiparameter persistence is based on the matching distance [dAmico2010, BiCeFrGi11, CeDFFeFrLa13, KeLeOu20, Landi2022, BjKe23]. This pseudo-metric is introduced by observing that, given a continuous function $\boldsymbol{{\varphi}}:X\to{\mathbb{R}}^{n}$ representing the data of interest, we can consider monoparametric filtrations of the topological space $X$ associated with lines in ${\mathbb{R}}^{n}$ having direction given by a vector $\boldsymbol{{a}}=({a}_{1},\ldots,{a}_{n})$ such that $\sum_{i=1}^{n}a_{i}=1$ and $a_{1},\ldots,a_{n}\geq 0$. Each such line, parametrised as $\boldsymbol{b}+u\boldsymbol{{a}}$ with $\boldsymbol{b}=(b_{1},\ldots,b_{n})$ and $\sum_{i=1}^{n}b_{i}=0$, induces a filtration $\{X^{(\boldsymbol{a},\boldsymbol{b})}_{u}\}_{u\in{\mathbb{R}}}$ of $X$ defined by the subsets $X^{(\boldsymbol{a},\boldsymbol{b})}_{u}$ consisting of all points in $X$ whose image under $\boldsymbol{{\varphi}}$ has coordinates less than or equal to those of $\boldsymbol{b}+u\boldsymbol{{a}}$. From an analytical viewpoint, this construction is equivalent to considering, for each function $\boldsymbol{\varphi}=({\varphi}_{1},\ldots,{\varphi}_{n}):X\to{\mathbb{R}}^{n}$, the sublevel sets of the real-valued function
@@
In recent years, the extended Pareto Grid has been introduced to compute points in the persistence diagram of $\boldsymbol{{\varphi}}^{*}_{\boldsymbol{a},\boldsymbol{b}}$, and to characterise at which $(\boldsymbol{a},\boldsymbol{b})$ the matching distance is attained [coher_match, EFQT2023, FrMoQuTo25].
While the study of the matching distance is geometrically interesting and theoretically rich it involves 2(n1)2(n-1) independent parameters and depends non-smoothly on the function $\boldsymbol{{\varphi}}$, leading to computational challenges and monodromy [coher_match]. We address this issue by considering the operator that maps a function
@@
$\operatorname{cmd}_{k}(\boldsymbol{\varphi},\boldsymbol{\psi}):=\max_{\begin{subarray}{c}t_{1},\ldots,t_{n}\geq 0\\ t_{1}+\ldots+t_{n}=1\end{subarray}}d_{B}\!\left(\operatorname{Dgm}_{k}(\boldsymbol{\varphi}^{\boldsymbol{t}}),\operatorname{Dgm}_{k}(\boldsymbol{\psi}^{\boldsymbol{t}})\right).$
@@
In this paper, we focus on the case $n=2$. We prove that the convex matching distance is stable with respect to the uniform norm $\|\boldsymbol{\varphi}-\boldsymbol{\psi}\|_{\infty}$ (Proposition 2) and establish a Position Theorem (Theorem 2), analogous to the one already known for the matching distance [coher_match, FrMoQuTo25]. This theorem allows one to recover, from the Pareto grid of $\boldsymbol{\varphi}$, the coordinates of the points in the persistence diagram of $\boldsymbol{\varphi}^{t}:=(1-t){\varphi}_{1}+t{\varphi}_{2}$, for every $t\in[0,1]$. Finally, in Theorem 3, we show that the convex matching distance coincides with the maximal bottleneck distance between the persistence diagrams of the functions $\boldsymbol{\varphi}^{t}$ and $\boldsymbol{\psi}^{t}$ for $t$ belonging to a suitable set of special values, thus paving the way for an efficient method to compute the convex matching distance.
The proposed strategy can be reinterpreted within the framework of Group Equivariant Non-Expansive Operators (GENEOs). Our new pseudo-metric is indeed associated with the GENEO $F_{t}$ that maps each function $\boldsymbol{\varphi}$ to the function $\boldsymbol{\varphi}^{t}$. GENEOs have been employed in TDA to tune the invariance properties of persistence diagrams to the requirements of specific applications [FrJa16], and in geometric and topological deep learning to design efficient neural architectures [bergomi2019towards, BoBoBrFrQu23, BoFeFr25, BoFrMi25, LaMiBoFrSo25, CoBoGiGiPeFr25].
Interestingly, infinitely many other examples of parametric families of GENEOs can be defined. For instance:
@@
Each of these operators could be used to define new pseudo-metrics in TDA.
The approach described in this paper is related to, but distinct from, the persistent homology transform (PHT) [TuMuBo14]. The PHT is a topological transform that takes as input a subset of a Euclidean space and associates to every unit vector the collection of persistence diagrams of the height functions on that subset in the direction given by the vector. A distance between two subsets is then defined by integrating over the sphere the distances between the respective persistence diagrams. The convex matching distance, instead, is obtained as a maximum over the unit vectors of non-negative slope, thus, over a smaller set. Considering only vectors of positive slope is a crucial aspect of our techniques because, together with some regularity assumptions, it enables the use of a convenient differential geometry tool, called the Pareto grid. Another key difference between the two approaches lies in their scope. While the PHT is designed to recover the geometry of the object $X$, our method, based on convex combinations of a vector-valued function, focuses on making inferences about the function whose domain is $X$.
@@ 2 Preliminaries
In this section, we recall the bottleneck distance in monoparameter persistence and the matching distance in biparameter persistence. To do that, some preliminary notions about persistence theory is needed, and certain hypothesis need to be assumed. However, as this is fairly standard, we will not delve into details, and refer the reader to [FrMoQuTo25, Section 2 and Appendix A] for a compact treatment of preliminaries. The following is a complete list of symbols used in this article  [^1] Note that these symbols might differ slightly from those used in [FrMoQuTo25].:
@@
Let ${\varphi},\psi\colon X\to{\mathbb{R}}$ be continuous functions. The $k$-bottleneck distance [Cohen-Steiner2007] is a metric between persistence diagrams defined by
@@
Let $\boldsymbol{{\varphi}},\boldsymbol{\psi}\colon X\to{\mathbb{R}}^{2}$ be continuous functions. The $k$-th matching distance [CeDFFeFrLa13] is a metric between the persistence Betti number functions (rank invariants) defined by
@@ 3 Convex matching distance
It follows that the map $t\mapsto\boldsymbol{\varphi}^{t}$ is $\lVert\varphi_{1}-\varphi_{2}\rVert_{\infty}$-Lipschitz. From the stability of the matching distance [CeDFFeFrLa13]—which implies $d_{B,k}(\boldsymbol{\varphi}^{t_{1}},\boldsymbol{\varphi}^{t_{2}})\leq\lVert\boldsymbol{\varphi}^{t_{1}}-\boldsymbol{\varphi}^{t_{2}}\rVert_{\infty}$ and $d_{B,k}(\boldsymbol{\psi}^{t_{1}},\boldsymbol{\psi}^{t_{2}})\leq\lVert\boldsymbol{\psi}^{t_{1}}-\boldsymbol{\psi}^{t_{2}}\rVert_{\infty}$—and by applying the triangle inequality, we get
@@
If $\boldsymbol{{\varphi}}$ is $k$-times differentiable, with $k$ possibly infinite, then so it is the function $\boldsymbol{{\varphi}}^{t}$ for each $t\in[0,1]$, due to being defined as the composition of $\boldsymbol{{\varphi}}=({\varphi}_{1},{\varphi}_{2})$ with the smooth function $(s_{1},s_{2})\mapsto(1-t)s_{1}+ts_{2}$. This is a key difference between $\operatorname{\mathbf{cmd}}$ and the classical matching distance, for the operator sending $\boldsymbol{{\varphi}}$ to $\boldsymbol{{\varphi}}_{(\boldsymbol{a},\boldsymbol{b})}$ preserves continuity only.
@@
For any $t\in[0,1]$, by the stability of the bottleneck distance [Cohen-Steiner2007] and Remark 2,
@@ 4 Comparison with the matching distance
In this section, we compare the newly introduced convex matching distance with the classical matching distance [CeDFFeFrLa13].
@@ 5 Pareto grid and Position Theorem
We call the closures of the images of the connected components of $\mathbb{J}_{P}(\boldsymbol{{\varphi}})\setminus\mathbb{J}_{C}(\boldsymbol{{\varphi}})$ (proper) contours [coher_match], and denote by $\mathrm{Ctr}(\boldsymbol{{\varphi}})$ the set of contours of $\varphi$}. Given two smooth functions $\boldsymbol{{\varphi}},\boldsymbol{{\psi}}\colon M\to\mathbb{R}^{2}$, we set $\mathrm{Ctr}(\boldsymbol{{\varphi}},\boldsymbol{{\psi}}):=\mathrm{Ctr}(\boldsymbol{{\varphi}})\cup\mathrm{Ctr}(\boldsymbol{{\psi}})$. It is important to observe that Assumption 1 is generic within the class of smooth functions from $M$ to ${\mathbb{R}}^{2}$ [wan].
@@
Before presenting the main result of this section, we recall the following well-established theorem, which provides a useful relation between the persistence diagram of a smooth function and its critical values. This statement is a technical but standard fact in the literature, and we include it here without proof (cf., e.g., Theorem 3.1 in [Simms_1964], and [Fr96]).
@@
We are now ready to state the main result of this section. Relying on Theorem 1, it provides a geometric characterisation of the points in the persistence diagram of $\boldsymbol{{\varphi}}^{t}$ in terms of the contours associated with $\boldsymbol{{\varphi}}$. Next theorem is analogous to [coher_match, Theorem 2]. In what follows, the symbol $\cdot$ denotes the dot product.
@@ 6 Special values
The intuition behind Definition 5.2 is the following: if $t\in[0,1]$ satisfies Definition 5.2, then $t$ is a value at which the optimal matching between $\operatorname{Dgm}(\boldsymbol{{\varphi}}^{t})$ and $\operatorname{Dgm}(\boldsymbol{{\psi}}^{t})$ may change abruptly. In particular, if $\sigma$ is such an optimal matching and $p_{1},p_{2},p_{3},p_{4}\in\operatorname{Dgm}(\boldsymbol{{\varphi}}^{t})\cup\operatorname{Dgm}(\boldsymbol{{\psi}}^{t})$ are such that $\operatorname{cost}\sigma=w_{1}-w_{2}=c(w_{3}-w_{4})$, where $w_{i}$ is a finite coordinate of $p_{i}$, $i=1,\ldots,4$, then, in a neighbourhood of $t$, the optimal matching could be either the one matching $p_{1}$ to $p_{2}$ or the one matching $p_{3}$ to $p_{4}$, but not the other. A similar motivation led to the definition of special set in [EFQT2023], relatively to the classical matching distance.

### v2

@@ Abstract
We introduce the convex matching distance, a novel metric for comparing functions with values in the real plane. This metric measures the maximal bottleneck distance between the persistence diagrams associated with the convex combinations of the two function components. Similarly to the traditional matching distance, the convex matching distance aggregates the information provided by two real-valued components. However, whereas the matching distance depends on two parameters, the convex matching distance depends on only one, offering improved computational efficiency. We further show that the convex matching distance can be more discriminative than the traditional matching distance in certain cases, although the two metrics are generally not comparable. Moreover, we prove that the convex matching distance is stable and characterize the coefficients of the convex combination at which it is attained. Finally, we demonstrate that this new aggregation framework benefits from the computational advantages provided by the Pareto grid, a collection of curves in the plane whose points lie in the image of the Pareto critical set associated with functions assuming values on the real plane. Experimental validation on MNIST digits, synthetic shapes, and chaotic attractors suggests that the convex matching distance provides a reliable and efficient alternative to the matching distance, at a significantly lower computational cost.
@@ 1 Introduction
In recent years, research on multiparameter persistent homology has grown substantially, driven by the need to apply Topological Data Analysis (TDA) techniques to data represented by multifiltrations, as occurs, for example, when analysing or comparing data described by vector-valued functions [31, 7, 17, 14, 16, 15, 18, 23, 19, 22, 38, 3, 34]. However, a fully developed theoretical framework enabling efficient applications of TDA in this setting is still lacking. A widely adopted comparison method in multiparameter persistence is based on the matching distance [d’Amico2010, 8, 20, 36, 4, 9]. This pseudo-metric is introduced by observing that, given a continuous function $\boldsymbol{{\varphi}}:X\to{\mathbb{R}}^{n}$ representing the data of interest, we can consider monoparametric filtrations of the topological space $X$ associated with lines in ${\mathbb{R}}^{n}$ having direction given by a vector $\boldsymbol{{a}}=({a}_{1},\ldots,{a}_{n})$ such that $\sum_{i=1}^{n}a_{i}=1$ and $a_{1},\ldots,a_{n}>0$. Each such line, parametrised as $\boldsymbol{b}+u\boldsymbol{{a}}$ with $\boldsymbol{b}=(b_{1},\ldots,b_{n})$ and $\sum_{i=1}^{n}b_{i}=0$, induces a filtration $\{X^{(\boldsymbol{a},\boldsymbol{b})}_{u}\}_{u\in{\mathbb{R}}}$ of $X$ defined by the subsets $X^{(\boldsymbol{a},\boldsymbol{b})}_{u}$ consisting of all points in $X$ whose image under $\boldsymbol{{\varphi}}$ has coordinates less than or equal to those of $\boldsymbol{b}+u\boldsymbol{{a}}$. From an analytical viewpoint, this construction is equivalent to considering, for each function $\boldsymbol{\varphi}=({\varphi}_{1},\ldots,{\varphi}_{n}):X\to{\mathbb{R}}^{n}$, the sublevel sets of the real-valued function
@@
In recent years, the extended Pareto Grid has been introduced to compute points in the persistence diagram of $\boldsymbol{{\varphi}}^{*}_{\boldsymbol{a},\boldsymbol{b}}$, and to characterise at which $(\boldsymbol{a},\boldsymbol{b})$ the matching distance is attained [21, 26, 30].
While the study of the matching distance is geometrically interesting and theoretically rich it involves 2(n1)2(n-1) independent parameters and depends non-smoothly on the function $\boldsymbol{{\varphi}}$, leading to computational challenges and monodromy [21]. We address this issue by considering the operator that maps a function
@@
$\operatorname{\mathbf{cmd}}_{k}(\boldsymbol{\varphi},\boldsymbol{\psi}):=\max_{\begin{subarray}{c}t_{1},\ldots,t_{n}\geq 0\\ t_{1}+\ldots+t_{n}=1\end{subarray}}d_{B}\!\left(\operatorname{Dgm}_{k}(\boldsymbol{\varphi}^{\boldsymbol{t}}),\operatorname{Dgm}_{k}(\boldsymbol{\psi}^{\boldsymbol{t}})\right).$
@@
The theoretical framework from which the convex matching distance derives is that of group equivariant non-expansive operators (GENEOs), which were introduced ten years ago to restrict the invariance of persistence diagrams in Topological Data Analysis [29]. In particular, the convex matching distance originates from the paper [5], where it was proved that selecting an arbitrary family of GENEOs leads to the construction of new stable matching distances (Proposition 9 and Theorem 10; see also [28]). In the present paper, we focus on the family of operators $F_{\boldsymbol{t}}(\boldsymbol{{\varphi}})=\boldsymbol{{\varphi}}^{\boldsymbol{t}}$, which yields the convex matching distance, owing to the interesting computational properties of this metric. Interestingly, infinitely many other examples of parametric families of GENEOs can be defined. For instance, restricting ourselves to monoparametric families, we may consider the following operators:
@@
Each of these operators could be used to define new pseudo-metrics in TDA. We emphasize that the theory of GENEOs also has interesting applications beyond classical Topological Data Analysis (cf., e.g., [5, 40, 1, 10, 12, 25, 27, 11, 13, 37, 2]).
For alternative approaches to this idea of a unifying theory to generate distances in persistence, we refer to [6] where projected distances between multiparameter persistence modules are introduced in the context of sheaf theoretic persistence. The convex matching distance can also be framed in this context.
The approach described in this paper is related to, but distinct from, the persistent homology transform (PHT) [44]. The PHT is a topological transform that takes as input a subset of a Euclidean space and associates to every unit vector the collection of persistence diagrams of the height functions on that subset in the direction given by the vector. A distance between two subsets is then defined by integrating over the sphere the distances between the respective persistence diagrams. The convex matching distance, instead, is obtained as a maximum over the unit vectors of non-negative slope, thus, over a smaller set. Considering only vectors of positive slope is a crucial aspect of our techniques because, together with some regularity assumptions, it enables the use of a convenient differential geometry tool, called the Pareto grid. Another key difference between the two approaches lies in their scope. While the PHT is designed to recover the geometry of the object $X$, our method, based on convex combinations of a vector-valued function, focuses on making inferences about the function whose domain is $X$.
In this paper, we study the convex matching distance in the case $n=2$. We prove that the convex matching distance is stable with respect to the uniform norm $\|\boldsymbol{\varphi}-\boldsymbol{\psi}\|_{\infty}$ (Proposition 2) and establish a Position Theorem (Theorem 2), analogous to the one already known for the matching distance [21, 30]. This theorem allows one to recover, from the Pareto grid of $\boldsymbol{\varphi}$, the coordinates of the points in the persistence diagram of $\boldsymbol{\varphi}^{t}:=(1-t){\varphi}_{1}+t{\varphi}_{2}$, for every $t\in[0,1]$. Finally, in Theorem 3, we show that the convex matching distance coincides with the maximal bottleneck distance between the persistence diagrams of the functions $\boldsymbol{\varphi}^{t}$ and $\boldsymbol{\psi}^{t}$ for $t$ belonging to a suitable set of special values, thus paving the way for an efficient method to compute the convex matching distance.
We validate the practical effectiveness of the convex matching distance through experiments on three benchmark datasets: MNIST digit classification, synthetic 3D shapes under varying noise levels, and chaotic dynamical attractors. Our results highlight that the convex matching distance achieves comparable classification performance to the matching distance computed on a fine $11\times 11$ parameter grid, while requiring only $11$ persistence diagram computations instead of $121$. Moreover, we observe that the convex matching distance can outperform coarser approximations of the matching distance, confirming that the convex parametrization captures the essential discriminative information of the bifiltration. These findings establish the convex matching distance as a computationally efficient and theoretically grounded alternative to existing multiparameter persistence metrics.
@@ 2 Preliminaries
In this section, we recall the bottleneck distance in monoparameter persistence and the matching distance in biparameter persistence. To do that, some preliminary notions about persistence theory is needed, and certain hypothesis need to be assumed. However, as this is fairly standard, we will not delve into details, and refer the reader to [30, Section 2 and Appendix A] for a compact treatment of preliminaries. The following is a complete list of symbols used in this article  [^1] Note that these symbols might differ slightly from those used in [30].:
@@
Let ${\varphi},\psi\colon X\to{\mathbb{R}}$ be continuous functions. The $k$-bottleneck distance [24] is a metric between persistence diagrams defined by
@@
Let $\boldsymbol{{\varphi}},\boldsymbol{\psi}\colon X\to{\mathbb{R}}^{2}$ be continuous functions. The $k$-th matching distance [20] is a metric between the persistence Betti number functions (rank invariants) defined by
@@ 3 Convex matching distance
It follows that the map $t\mapsto\boldsymbol{\varphi}^{t}$ is $\lVert\varphi_{1}-\varphi_{2}\rVert_{\infty}$-Lipschitz (see also [6, Proposition 7.9]). From the stability of the matching distance [20]—which implies $d_{B,k}(\boldsymbol{\varphi}^{t_{1}},\boldsymbol{\varphi}^{t_{2}})\leq\lVert\boldsymbol{\varphi}^{t_{1}}-\boldsymbol{\varphi}^{t_{2}}\rVert_{\infty}$ and $d_{B,k}(\boldsymbol{\psi}^{t_{1}},\boldsymbol{\psi}^{t_{2}})\leq\lVert\boldsymbol{\psi}^{t_{1}}-\boldsymbol{\psi}^{t_{2}}\rVert_{\infty}$—and by applying the triangle inequality, we get
@@
If $\boldsymbol{{\varphi}}$ is $k$-times differentiable, with $k$ possibly infinite, then so it is the function $\boldsymbol{{\varphi}}^{t}$ for each $t\in[0,1]$, due to being defined as the composition of $\boldsymbol{{\varphi}}=({\varphi}_{1},{\varphi}_{2})$ with the smooth function $(s_{1},s_{2})\mapsto(1-t)s_{1}+ts_{2}$. This is a key difference between the convex matching distance and the classical matching distance, for the operator sending $\boldsymbol{{\varphi}}$ to $\boldsymbol{{\varphi}}^{*}_{\boldsymbol{a},\boldsymbol{b}}$ preserves continuity only.
@@
For any $t\in[0,1]$, by the stability of the bottleneck distance [24] and Remark 2,
@@ 4 Comparison with the matching distance
In this section, we compare the newly introduced convex matching distance with the classical matching distance [20].
@@ 5 Pareto grid and Position Theorem
We call the closures of the images of the connected components of $\mathbb{J}_{P}(\boldsymbol{{\varphi}})\setminus\mathbb{J}_{C}(\boldsymbol{{\varphi}})$ (proper) contours [21], and denote by $\mathrm{Ctr}(\boldsymbol{{\varphi}})$ the set of contours of $\varphi$}. Given two smooth functions $\boldsymbol{{\varphi}},\boldsymbol{{\psi}}\colon M\to\mathbb{R}^{2}$, we set $\mathrm{Ctr}(\boldsymbol{{\varphi}},\boldsymbol{{\psi}}):=\mathrm{Ctr}(\boldsymbol{{\varphi}})\cup\mathrm{Ctr}(\boldsymbol{{\psi}})$. It is important to observe that Assumption 1 is generic within the class of smooth functions from $M$ to ${\mathbb{R}}^{2}$ [46].
@@
Before presenting the main result of this section, we recall the following well-established theorem, which provides a useful relation between the persistence diagram of a smooth function and its critical values. This statement is a technical but standard fact in the literature, and we include it here without proof (cf., e.g., Theorem 3.1 in [41], and [32]).
@@
We are now ready to state the main result of this section. Relying on Theorem 1, it provides a geometric characterisation of the points in the persistence diagram of $\boldsymbol{{\varphi}}^{t}$ in terms of the contours associated with $\boldsymbol{{\varphi}}$. Next theorem is analogous to [21, Theorem 2]. In what follows, the symbol $\cdot$ denotes the dot product.
@@ 6 Special values
The intuition behind Definition 5.2 is the following: if $t\in[0,1]$ satisfies Definition 5.2, then $t$ is a value at which the optimal matching between $\operatorname{Dgm}(\boldsymbol{{\varphi}}^{t})$ and $\operatorname{Dgm}(\boldsymbol{{\psi}}^{t})$ may change abruptly. In particular, if $\sigma$ is such an optimal matching and $p_{1},p_{2},p_{3},p_{4}\in\operatorname{Dgm}(\boldsymbol{{\varphi}}^{t})\cup\operatorname{Dgm}(\boldsymbol{{\psi}}^{t})$ are such that $\operatorname{cost}\sigma=w_{1}-w_{2}=c(w_{3}-w_{4})$, where $w_{i}$ is a finite coordinate of $p_{i}$, $i=1,\ldots,4$, then, in a neighbourhood of $t$, the optimal matching could be either the one matching $p_{1}$ to $p_{2}$ or the one matching $p_{3}$ to $p_{4}$, but not the other. A similar motivation led to the definition of special set in [26], relatively to the classical matching distance.